OpenStudy (anonymous):

If f(x)=7-xsquared f(4a)=?

OpenStudy (anonymous):

f(x)=7-x^2 f(4a) = 7-(4a)^2 = 7-16a^2

OpenStudy (anonymous):

oh, i get it thanks

OpenStudy (anonymous):

welcome